1. A large tank is filled to capacity with 1000 L of pure water. Brine containing 0.25 kg of salt per liter is pumped into the tank at a rate of 10 L/min. The well-mixed solution is pumped out at the same rate. Find the number A(t) of kilograms of salt in the tank at time t.

A(t) = ___ kg

2. A tank contains 250 liters of fluid in which 10 grams of salt is dissolved. Brine containing 1 gram of salt per liter is then pumped into the tank at a rate of 5 L/min; the well-mixed solution is pumped out at the same rate. Find the number A(t) of grams of salt in the tank at time t.

A(t) = ___
1
Expert's answer
2020-06-22T17:43:33-0400

The process is described by differential equation "dA=\frac{mt-A}{V}Mdt" where "M" -is the brine of salt rate "(L\/min)" , "m" - is salt rate "(kg\/min)" , "V" - is initial volume of water in a tank "(L)" .The solution to this equation is "A(t)=m(t-\frac{V}{M}(1-e^{-\frac{Mt}{V}})+Ce^{-\frac{Mt}{V}})" then

  1. As "A(0)=0" hence "C=0" then "A(t)=2.5(t-100(1-e^{-0.01t}))kg"
  2. As "A(0)=10g" hence "C=2min" then "A(t)=5(t-50(1-e^{-0.02t})+2e^{-0.02t})g"